Circular No. 6905

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998bu IN NGC 3368 K. Ayani, Bisei Astronomical Observatory (BAO); H. Nakatani, Fuchu,

Hiroshima, Japan; and H. Yamaoka, Kyushu University, report that a wideband spectrogram (range 400–750 nm) of SN 1998bu (IAUC 6899, 6902) was obtained with the BAO 1.01-m telescope on May 14.5 UT. Preliminary reduction reveals that it has deep and broad Si II absorption at 612 nm (at the rest frame of NGC 3368), which indicates that SN 1998bu is of type Ia. Other broad absorption lines are seen at 575, 544, 527, 486, and 427 nm. The expansion velocity derived from the shift of the Si II absorption minimun is ∼ 11 000 km/s.

P. Meikle, M. Hernandez, and A. Fassia, Imperial College, London, report: “A spectrum of SN 1998bu was obtained by J. Iglesias with the William Herschel Telescope of the Isaac Newton Group, La Palma, on May 12.9 UT. The ISIS spectrograph (range 315–898 nm, resolution 0.8 nm) yields a spectrum of a type-Ia supernova a few days before maximum light. It is similar to the spectrum of SN 1990N at one week before B maximum (Leibundgut et al. 1991, Ap.J. 371, L23). The Si II absorption minimum is at 611.5 nm (host galaxy rest frame, adopting the heliocentric radial velocity +744.8 km/s for M96 given by Munari et al. on IAUC 6902).

B. A. Skiff, Lowell Observatory, reports that C. Faranda, Jr., Lauderhill, FL, obtained a frame of NGC 3368 on May 3.14 UT (0.25-m reflector + unfiltered CCD) that shows a prediscovery image of SN 1998bu at mag 16.5 ± 0.3 (using a V -band comparison-star sequence).

Visual magnitude estimates: May 13.22 UT, 12.1 (C. E. Spratt, Victoria, BC); 13.86, 12.2 (A. Baransky, Kiev, Ukraine); 13.90, 11.8 (K. Hornoch, Lelekovice, Czech Republic).

XTE J1806–246 R. I. Hynes, P. Roche, and C. A. Haswell, Sussex University, report:

“Images taken on May 6.2 UT with the Jacobus Kapteyn Telescope at La Palma of the field of XTE J1806–246 (IAUC 6891) show a star coincident with the suggested radio counterpart (IAUC 6900). It is close to the limit of detectability in V, R, and I images (estimated at V ∼ 22) at α = 18^{h}06^{m}50^{s}.75, δ = −24^{o}35′28′′.6 (equinox 2000.0; ± 1′′.0). This is the only object coincident with the radio source, and we suggest that it may be the optical counterpart to XTE J1806–246.”

1998 May 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green